Exercise 2. Suppose a researcher wants to estimate the mean weight of all Karate players in a tournament. He took a random sample of 30 volleybal players and ther height (in centimeters). Their weight was recorded as follows: 55 45 45 55 60 59 57 66 80 70 64 65 67 65 65 65 82 73 50 55 75 66 77 65 82 71 65 52 75 What is the mean height of the 30 volleyball players in the collegiate tournament?​

1
Expert's answer
2022-04-21T06:38:46-0400

"\mu =\frac{\sum x}{29}=(55\times 3+45\times2+60+59+57+66\times2+80+70+64+65\times6+67+82\times2+73+50+75\times2+77+71+52)/29=1871/29=64.5"